OpenStudy (anonymous):

n(n ‒ 3) + 2(n ‒ 3) = ( )(n ‒ 3) That video didn't help me at all.

OpenStudy (baru):

i don't think we can explain it better than the video. I think this has to be explained in person, ask your teachers or your parents :)

OpenStudy (anonymous):

I don't have a teacher.

OpenStudy (baru):

:( where did you get this question then?

OpenStudy (anonymous):

Plato and my dad won't help me either.

OpenStudy (baru):

so you know PEMDAS?

OpenStudy (anonymous):

Yes. I don't under stand the pulling the letters out... Why would I do that?!

OpenStudy (baru):

ah..ok firstly, they are not letters, they are letters which represent 'numbers' and the pulling them out part..its just property of numbers, its a law of mathematics do you know what 'factors' are?

OpenStudy (anonymous):

No

OpenStudy (baru):

well....you should have learn what factors are by now, judging by the question you posted. there are a few simple concepts which you should be knowing in order to answer that question, some of them are representing numbers as letters factors distributive property of multiplication over addition Khan academy is a good place to start

OpenStudy (anonymous):

Whatever I just I quit.

OpenStudy (baru):

well, thats upto you, but this is very basic math which you will actually need in life, even if you want nothing to do with math or science

OpenStudy (anonymous):

I'm sure I won't need math to drive.

OpenStudy (revircs):

Just plug in the leading coefficients from the left side. n(n-3)+2(n-3)=( ? )(n-3) n(n-3)+2(n-3)=(n+2)(n-3)

OpenStudy (revircs):

The further explanation would be to distribute the leading coefficients and factoring from there.